# The Translation Trap Hypothesis for Multilingual Generation with Large Language Models The Translation Trap Hypothesis: Translation as a Constraint on Multilingual Generation

### v1

@@ Abstract
Multilingual generation with large language models (LLMs) is often of poor quality for mid- to low-resource languages. Building on insights from interpretability, we demonstrate the existence of an implicit task-solvingtranslation pipeline for generation, whereby the model first solves the required task in a largely target-language-agnostic manner, and subsequently translates answer concepts into the intended target language. We hypothesize that the failure of the translation stage is an important culprit for the observed low quality of final outputs, and formalize this as the translation barrier hypothesis. We test this hypothesis for a word translation task across language pairs, using logit lens to observe model processing in intermediate layers. We find that a significant portion of overall failures indeed stems from translation failure, or the model’s inability to translate correctly solved intermediate concepts into the target language. This is especially true for low-resource target languages. Our results highlight an important hurdle for end-to-end multilingual generation, and lend guiding insights for future work seeking to improve multilinguality in LLMs.111github.com/JHU-CLSP/translation-barrier
@@ 1 Introduction
Work from the field of interpretability shows that intermediate layer representations of multilingual LLMs are close to English representations even when prompted to generate in other languages, and that the models demonstrate language-specific behavior only in late layers of the model (Wendler et al., 2024; Foroutan et al., 2022; Tang et al., 2024; Kojima et al., 2024). Building upon this intuition, we posit a model-internal task-solvingtranslation cascade, with the middle layers of the LLM responsible for task-solving, or discovering the required output concepts, in a target-language-agnostic manner, and the last few layers responsible for realizing those concepts in the target language, through translation.
We present the translation barrier hypothesis, illustrated and stated in Figure 1, regarding the poor quality of multilingual generation. This hypothesis has important consequences for end-to-end multilingual generation. If translation failure, as opposed to task-solving failure, is the dominant culprit for the generation gap, then we can modularize or investigate architectural alternatives for the final LLM layers, or seek to induce better transfer in other ways, while still relying on powerful LLM task processing for end-to-end generation in LRLs. On the other hand, if task-solving capabilities themselves fail for LRL targets, explicit cascading methods become more viable options, calling for more investment in external specialized HRL-LRL MT systems.
Our work explores the translation barrier on a word translation task, for two mainstream multilingual models. We choose word translation as a minimal generation task: in this simple case, task-solving consists simply of realizing the semantics or the content of the input term, whereas translation refers to generating the target language term for that concept. We use logit lens to retrieve model-internal hypotheses across different layers, assessing whether the model has arrived at the appropriate semantics for a particular input regardless of language; we also assess final on-target accuracy at the final layer. This allows us to demonstrate the existence of the posited pipeline. We then test our hypothesis, by quantifying the extent to which either stage in the pipeline is responsible for final failure. We make the following contributions:
We introduce the translation barrier hypothesis, highlighting a key mechanism of failure for multilingual generation with LLMs.
Working with a word translation task, we provide evidence for the existence of an implicit task-solvingtranslation pipeline for LLM generation over (source) (target) language pairs, developing a framework within which we can assess and compare the contribution of task-solving and translation failure to total final failure.
@@
We find that translation failure dominates () final failure for and of all language pairs for aya-23 and llama-3.1 respectively for our task, and that it especially forms the bottleneck for generation in low-resource or unsupported target languages. On the other hand, we find that task-solving is the bottleneck in the case of a low-resource source language.
@@ 2 Experimental Setup
Given a word translation task between a source and target language, we are interested in judging the semantics of intermediate layer outputs regardless of language. We create a multiparallel dataset consisting of translations of words aligned across all languages of interest. We do this by taking source words in English, split equally among nouns, verbs, adjectives, and adverbs, and using Google Translate API222translate.google.com to collect a set of translations per word, including synonyms, for all languages and for English itself. This allows us to construct a lexicon between any source and target language for the same concepts, as well as assess off-target accuracy in all other languages per concept. We manually inspected the source words and lexicons for correctness and coverage (see Appendix B for more details).
@@
By virtue of their data-hungry nature, mainstream LLMs only support a limited set of usually high- and mid-resource languages, but may be used and are claimed to demonstrate incident multilingualism in other languages (Blevins and Zettlemoyer, 2022). We study target languages, including all languages supported by either model,333We count simplified and traditional Chinese as separate; thus we say that aya-23 supports languages. Thai is the only language supported by llama-3.1 but not aya-23. and a mix of additional mid- and low-resource languages unsupported by both models. See Appendix A for the full list of supported and unsupported languages. We use three different source languages against all the above targets: a high-resource supported language (Spanish), a mid-resource supported language (Hindi), and a mid-to-low resource unsupported language (Telugu). Note that “source” language refers to that of the translation task input; the task instruction is always provided in English. We consider all the above languages as candidates for task-solving languages in intermediate model layers.
@@ 3 Quantifying Task-solving and Translation failure
The translation barrier hypothesis presupposes the existence of a model-internal task-solvingtranslation pipeline (validated and discussed in § 4), with model success predicated on task-solving success followed by translation success. We can study this hypothesis in the context of a given task by quantifying the extent to which translation failure is responsible for poor final outputs, as opposed to task-solving failure.
@@
We use an interpretability tool called logit lens (Nostalgebraist, 2020) to obtain text outputs from intermediate layers. Let be an input prompt and be the output of layer . Logit lens applies the unembedding matrix and softmax to to produce a probability distribution over the vocabulary at layer :
@@
For each layer output, we are interested in its task-solving accuracy, measured by the exact match metric , and its language. We use our dataset language tag for the layer output in case of a match against our listed language equivalents. For inaccurate outputs, we perform language ID on the layer output with NLLB language ID model (NLLB Team et al., 2022).444huggingface.co/facebook/fasttext-language-identification/ We only retain the tag if it belongs to our set of target languages, which includes all languages supported by the model i.e. plausible intermediate languages, and discard it if not. This is because intermediate layer output text may be gibberish or malformed, in which case we expect the language ID model to return a noisy random tag. See Appendix E for a manual evaluation of the reliability of the resulting tags. We track accuracy and language of output for the last ten layers of each model.555We inspected all layers with preliminary experiments and found that intermediate accuracy emerges in a model- and language- dependent manner around this layer. Including earlier intermediate layers yielded similar results for our experiments.
@@ 4.2 Characterizing task-solving languages
We use the proportion of the target language among all accurate layer outputs to track the switch between the two stages, since this will be high in the translation stage, and plot this by layer for five representative languages in Figure 3 for both models.666We also look at the proportion of target language outputs regardless of accuracy by layer; this results in similar trends. See Appendix H for complete plots.
In general, both models show a significance increase in target language presence around the last four layers. We imagine that this is model- and task- dependent. We also compute the specific layer with the maximum increase in target language presence from the previous layer ranges from layer, and find that this ranges to over different language pairs (see Appendix G for details). We found no correlation between the layer of switch and accuracy at the final layer over all language pairs.
@@
As per the above, we consider all layers up to the fourth-last layer as task-solving layers,777This is a conservative but heuristic approximation: as shown in Figure 3, the switch between the stages is fairly fluid around the last few layers. and look at the distribution over languages for all identified intermediate correct (off-target) answers in these layers. We show the aggregate distribution over all language pairs in Figure 4 for both models.
@@ 4.3 Quantifying the translation barrier
Next, we look at intermediate accuracy, final accuracy, and by language pair and model to understand how these vary in more detail. We show these results for target languages against each source language for both models in Figure 5, with the complete results for all target languages in Appendix J. Broadly, high intermediate accuracy indicates successful task-solving; if it is accompanied by high final accuracy, that indicates a subsequently successful translation stage. As discussed in § 3, is a function of these quantities, measuring the proportion of failure cases (i.e. the complement of final accuracy) that can be attributed to translation loss (representing cases that were correctly task-solved but incorrectly translated).
@@
We find that translation failure dominates total failure () for and of all language pairs for and aya-23 and llama-3.1 respectively. This provides direct evidence for the translation barrier hypothesis for individual language pairs.
We further find distinct patterns for high-resource versus low-resource target languages. We observe that intermediate accuracy stays high even for low-resource unsupported target languages (like Nepali and Hebrew) for Hindi and Spanish as sources: in line with our previous observations, this indicates that the model is able to correctly task-solve even for low-resource targets. In contrast, and as expected, we observe that final accuracy is high for supported target HRLs (like German, Italian, and Portuguese, for both models) but drops drastically for low-resource target languages (like Tamil and Swahili), some of which are supported, such as for Greek with aya-23. High intermediate and final accuracy as in the first case (HRLs) means that is relatively low: a larger part of total failure comes from task-solving errors, and the model is more capable of generating successfully in the case of task-solving success. However, high intermediate accuracy and low final accuracy as in the second case (LRLs) indicates that translation loss accounts for a much bigger percentage of total error, going up to for spa-cat with aya-23, and for spa-mar with llama-3.1.
Note that with the exception of language pairs with English as a target, or with Telugu as a source, translation failure dominates total failure () for and of language pairs for aya-23 and llama-3.1. English is a special case: given that it dominates the internal representations of these models, generation into English may have a minimal or absent translation stage, meaning that task-solving success largely determines final success. We discuss the case of Telugu as a source below.
@@
The above discussion generally supports our understanding of target-language-agnostic task-solving stage followed by a target-language-specific translation stage. However, we note the caveat that there is a clear disparity between intermediate task accuracy across target languages, with the largest difference of percentage points between spa-eng and spa-amh with aya-23, as well as differences between supported languages ( percentage point difference between hin-eng and hin-jpn with aya-23). This indicates some level of entanglement between task-solving success and language of generation, meaning that the intended language of generation affects the model’s ability to task-solve to some extent. This hints that given the current state of LLMs, even in the hypothetical case of no translation loss, end-to-end processing for low-resource or unsupported languages would still fall behind that for HRLs due to imperfect transfer of task-solving abilities to LRLs (including some supported languages). This finding calls for an evaluation of the future of end-to-end multilingual generation with LLMs on a larger scale of languages, with one alternative being to invest in better explicit output-side LLM+MT cascades.
@@
We see that for both models, mean intermediate and final accuracy as well as mean are highest for Spanish, then Hindi, then Telugu (Table 1). It is unsurprising that intermediate accuracy and therefore final accuracy depend on the source language: task-solving success is naturally conditioned on comprehension of the task inputs.999While we do not experiment with different prompt languages other than English, we can imagine that the prompt language may also show a similar effect. These findings justify exploring an input-side MT+LLM cascade for different language task inputs even for supported languages.
This also explains why is lowest for Telugu: intuitively, with low intermediate accuracy, translation loss can only occur on very few inputs as a percentage of the total failure, which is dominated by failed task-solving cases. For example, since task-solving fails for of total inputs for Telugu-Cebuano with aya-23, task-solving failure dominates total failure even though the translation stage fails to convert nearly all correct intermediate outputs to correct final outputs.
4.4 Effect of scale
We repeat our experiments for language pairs on Llama-3.1-70B-Instruct and show the effect of using a larger model on in Figure 6. We find that the model shows similar intermediate accuracy on average as the model, but consistently higher final accuracy, or capability to produce the target language, resulting in lower . However, absolute values remain high. These preliminary results indicate that the translation barrier may continue to be a problem at scale. We invite future work in better understanding the translation barrier at different scales and for other tasks.
@@ 5 Related Work
Behavior of Multilingual LLMs
While multilingual LLMs are trained to perform well for several languages (Lin et al., 2022; Workshop et al., 2023; Aryabumi et al., 2024), their pre-training data is dominated by English (Xue et al., 2021; Chung et al., 2023; Li et al., 2025a). Etxaniz et al. (2024); Zhao et al. (2024) show that the explicit cascade, i.e. translating inputs into English before passing them into LLMs, often outperforms end-to-end inference. Further, works such as Wendler et al. (2024) and (Schut et al., 2025) apply logit lens to intermediate layers and observe that multilingual LLMs “think” in English, or that intermediate representation spaces of these models lie close to English. This is reinforced by works studying neuron activations for multilingual processing (Foroutan et al., 2022; Tang et al., 2024; Kojima et al., 2024), which find language-specific neurons only in later model layers. Our work draws from this intuition with the goal of characterizing the implicit pipeline for multilingual generation and diagnosing its failure.
Cross-lingual Representations
@@ 6 Conclusion
Our paper seeks to investigate the reasons for the failure of multilingual generation with LLMs. We track the language and accuracy of intermediate model layer outputs to demonstrate a task-solvingtranslation pipeline for multilingual generation. We introduce the translation barrier hypothesis, calling attention to translation failure as an important culprit for final failure, and validate it for a word translation task by comparing task-solving as well as final success over language pairs. Our findings raise important questions for future strategies aimed at improving multilinguality in LLMs.
@@
The main limitation of this work is that we only explore and validate the translation barrier hypothesis for a word translation task. It may be more difficult to assess task-solving success for more complex tasks such as sentence-level machine translation, summarization, or open-ended instruction following in a principled manner using artifacts such as intermediate layer representations, since the nature of correct intermediate answer concepts could be less clear to ascertain. There may also be additional complexities to take into account when extending our setup to these tasks: for example, whether task-solving occurs “globally” over many parts or subtasks of a single input or “locally” for different fragments of the input, or the manner in which the target language syntax interacts with the ordering of the concepts in the task-solving stage. Thus, our hypothesis may be difficult to verify for other tasks.
Further, we clarify that our findings regarding the role played by the translation barrier may be dependent on the nature and complexity of the task, as well as the usage of strategies such as task-specific finetuning or inference-time techniques such as in-context learning to boost multilingual performance.
We do not claim that the statement of the hypothesis holds for all tasks and settings; however, we believe that the line of investigation in our work has important implications for future directions to improve the multilingual capabilities in LLMs, specifically regarding the potential of end-to-end versus cascaded approaches for the same. Our goal is to invite discussion and further investigation, including potential counter-evidence from other tasks, regarding the most viable strategies for high quality generation in mid- to low-resource languages with LLMs.
@@
Further, we use English as our prompt language regardless of the source language (i.e. the language of the term to be translated); this is as per work that says that this performs better (Dey et al., 2024). Testing the effect of different prompt languages may add further nuance to our findings.
@@
We do not anticipate any negative consequences of this work, which is intended to diagnose poor multilingual capabilities in large language models. However, we acknowledge that different language communities may have various and diverse needs from language technologies that may diverge from the points of focus of mainstream NLP (Bird, 2020). All the models we used are publicly available for research purposes: Aya-23-8B has a CC-BY-NC license, and the Llama family of models have a Llama 3.1 Community License Agreement, permitting research use.

### v2

# The Translation Barrier Hypothesis: Multilingual Generation with Large Language Models Suffers from Implicit Translation Failure

@@ Abstract
Multilingual generation with large language models (LLMs) is often of poor quality for mid- to low-resource languages, but the causes for this are not well-understood. We first demonstrate the existence of an implicit task-solvingtranslation pipeline for generation, whereby the model first solves the required task in a largely target-language-agnostic manner, and subsequently translates answer concepts into the intended target language. We hypothesize that the failure of the translation stage, despite task-solving success, is an important culprit for the observed low quality of final outputs, and formalize this as the translation barrier hypothesis. We quantify the extent to which either stage in the pipeline is responsible for final failure for a word translation task across language pairs, and find that the translation barrier explains a dominant portion of error for a majority of language pairs, and is especially severe for low-resource target languages. Our results highlight an important bottleneck for end-to-end multilingual generation, relevant for future work seeking to improve multilinguality in LLMs. 111github.com/JHU-CLSP/translation-barrier
@@ 1 Introduction
Work from mechanistic interpretability shows that intermediate layer representations of multilingual LLMs are close to English representations even when prompted to generate in other languages, and that the models demonstrate language-specific behavior only in late layers of the model (Wendler et al., 2024; Foroutan et al., 2022; Tang et al., 2024; Kojima et al., 2024). Building upon this intuition, we posit a model-internal task-solvingtranslation cascade, with the middle layers of the LLM responsible for task-solving, or discovering the required output concepts, in a target-language-agnostic manner, and the last few layers responsible for realizing those concepts in the target language, through translation.
We present the translation barrier hypothesis, illustrated and stated in Figure 1, regarding the poor quality of multilingual generation. This hypothesis has important consequences for end-to-end multilingual generation. If translation failure, as opposed to task-solving failure, is the dominant culprit for low multilingual performance, then we can modularize or investigate architectural alternatives for the final LLM layers, or seek to induce better transfer in other ways, while still relying on powerful LLM task processing for end-to-end generation in LRLs. On the other hand, if task-solving capabilities themselves fail for LRL targets, explicit cascading methods become more viable options, calling for more investment in external specialized HRL-LRL MT systems.
Our work quantifies the translation barrier on a word translation task, for two mainstream multilingual models, in order to demonstrate the extent of this issue. We choose word translation as a minimal generation task: in this simple case, task-solving consists simply of realizing the semantics or the content of the input term, whereas translation refers to generating the target language term for that concept. We use logit lens to retrieve model-internal hypotheses across different layers, assessing whether the model has arrived at the appropriate semantics for a particular input regardless of language; we also assess the on-target accuracy of the final outputs. This allows us to demonstrate the existence of the posited pipeline. We then test our hypothesis, by quantifying the extent to which either stage in the pipeline is responsible for final failure.
We make the following contributions:
We formalize the translation barrier hypothesis, highlighting a key mechanism of failure for multilingual generation with LLMs.
Working with a word translation task, we provide evidence for the existence of an implicit task-solvingtranslation pipeline for LLM generation over (target) (source) language pairs, and develop a framework within which we can measure and compare the contribution of task-solving and translation failure to total final failure.
@@
We find that translation failure dominates () final failure for and of all language pairs for aya-23 and llama-3.1 respectively, and that it especially forms the bottleneck for generation in low-resource or unsupported target languages. On the other hand, we find that task-solving is the bottleneck in the case of a low-resource source language, and is also affected to some degree by the intended target language.
We provide case studies with a larger sized model and on a different task. These yield consistent findings with the above, inviting future work in understanding the translation barrier in different settings.
@@ 2 Experimental Setup
Given a word translation task between a source and target language, we are interested in judging the semantics of intermediate layer outputs regardless of language. We create a multiparallel dataset consisting of translations of words aligned across all languages of interest. We take source words in English, split equally among nouns, verbs, adjectives, and adverbs, and use the Google Translate API222translate.google.com to collect a set of translations per word, including synonyms, for all languages and for English itself. This allows us to construct a lexicon between any source and target language for the same concepts, as well as assess off-target accuracy in all other languages per concept. We manually inspected the source words and lexicons for correctness and coverage (see Appendix B for more details).
@@
By virtue of their data-hungry nature, mainstream LLMs only support a limited set of usually high- and mid-resource languages, but may be used and are claimed to demonstrate incident multilingualism in other languages (Blevins and Zettlemoyer, 2022). We study target languages, including all languages supported by either model,333We count simplified and traditional Chinese as separate; thus we say that aya-23 supports languages. Thai is the only language supported by llama-3.1 but not aya-23. and a mix of additional mid- and low-resource languages unsupported by both models. See Appendix A for the full list of supported and unsupported languages and their resource level. We use three different source languages against all the above targets: a high-resource supported language (Spanish), a mid-resource supported language (Hindi), and a mid-to-low resource unsupported language (Telugu). Note that “source” language refers to that of the translation task input; the task instruction is always provided in English. We consider all the above languages as candidates for task-solving languages in intermediate model layers.
@@ 3 Quantifying Task-solving and Translation failure
The translation barrier hypothesis presupposes the existence of a model-internal task-solvingtranslation pipeline (demonstrated and discussed in § 4), with model success predicated on task-solving success followed by translation success. We can study this hypothesis in the context of a given task by quantifying the extent to which translation failure is responsible for poor final outputs, as opposed to task-solving failure.
@@
We use logit lens (Nostalgebraist, 2020) to obtain text outputs from intermediate layers. Let be an input prompt and be the output of layer . Logit lens applies the unembedding matrix and softmax to to produce a probability distribution over the vocabulary at layer :
@@
For each layer output, we are interested in its task-solving accuracy, measured by the exact match metric , and its language. We use our dataset language tag for the layer output in case of a match against our listed language equivalents. For inaccurate outputs, we perform language ID on the layer output with the NLLB language ID model (NLLB Team et al., 2022).444huggingface.co/facebook/fasttext-language-identification/ We only retain the tag if it belongs to our set of target languages, which includes all languages supported by the model i.e. plausible intermediate languages, and discard it if not. This is because intermediate layer output text may be gibberish or malformed, in which case we expect the language ID model to return a noisy random tag. See Appendix E for a manual evaluation of the reliability of the resulting tags. We track accuracy and language of output for the last ten layers of each model.555Preliminary experiments show that including earlier intermediate layers does not affect measurements of task-solving accuracy. Logit lens is also unreliable for early layers (Belrose et al., 2023).
@@ 4.2 Characterizing task-solving languages
We use the proportion of the target language among all accurate layer outputs to track the switch between the two stages, since this will be high in the translation stage, and plot this for five languages in Figure 3 for both models.666See Appendix H for complete plots. We also look at the proportion of target language outputs regardless of accuracy by layer; this results in similar trends.
In general, both models show a significance increase in target language presence around the last four layers. We imagine that this is model- and task- dependent. We also compute the specific layer with the maximum increase in target language presence from the previous layer, and find that this ranges to over different language pairs (see Appendix G for details). We found no correlation between the layer of switch and accuracy at the final layer over all language pairs.
@@
As per the above, we consider all layers up to the fourth-last layer as task-solving layers,777This is a conservative heuristic approximation: as shown in Figure 3, the switch between the stages is fairly fluid around the last few layers. Results look similar for surrounding choices of layer. and look at the distribution over languages for all identified intermediate correct (off-target) answers in these layers. We show the aggregate distribution over all language pairs in Figure 4 for both models.
@@ 4.3 Quantifying the translation barrier
Next, we look at intermediate accuracy, final accuracy, and by language pair and model. We show these results for target languages against each source language for both models in Figure 5, with the complete results for all target languages in Appendix J. We find that translation failure dominates total failure () for and of all language pairs for and aya-23 and llama-3.1 respectively. This provides direct evidence for the translation barrier hypothesis for individual language pairs.
We further find distinct patterns for high-resource versus low-resource target languages. We observe that intermediate accuracy stays high even for low-resource unsupported target languages (like Nepali and Hebrew) for Hindi and Spanish as sources: in line with our previous observations, this indicates that the model is able to correctly task-solve even for low-resource targets. In contrast, and as expected, we observe that final accuracy is high for supported target HRLs (like German, Italian, and Portuguese, for both models) but drops drastically for low-resource target languages (like Tamil and Swahili), some of which are supported, such as for Greek with aya-23. Script and language family may naturally also contribute to this performance drop (e.g. llama-3.1 performs worse for Japanese and Arabic than for Romanian). High intermediate and final accuracy as in the first case (HRLs) means that is relatively low: a larger part of total failure comes from task-solving errors, and the model is more capable of generating successfully in the case of task-solving success. However, high intermediate accuracy and low final accuracy as in the second case (LRLs) indicates that translation loss accounts for a much bigger percentage of total error, going up to for spa-cat with aya-23, and for spa-mar with llama-3.1.
Leaving aside language pairs with English as a target, or with Telugu as a source, translation failure dominates total failure () for and of language pairs for aya-23 and llama-3.1. English is a special case: given that it dominates the internal representations of these models, generation into English may have a minimal or absent translation stage, meaning that task-solving success largely determines final success. We discuss the case of Telugu as a source below.
@@
The above discussion generally supports our understanding of target-language-agnostic task-solving stage followed by a target-language-specific translation stage. However, we note the caveat that there is a clear disparity between intermediate task-solving accuracy across target languages, with the largest difference of percentage points between spa-eng and spa-amh with aya-23, as well as differences between supported languages ( percentage point difference between hin-eng and hin-jpn with aya-23). This indicates some level of entanglement between task-solving success and language of generation, meaning that the intended language of generation affects the model’s ability to task-solve to some extent. This hints that even in the hypothetical case of no translation loss, end-to-end processing for low-resource or unsupported languages would still fall behind that for HRLs due to imperfect transfer of task-solving abilities to LRLs. This finding calls for an evaluation of the future of end-to-end multilingual generation with LLMs on a larger scale of languages, with one alternative being to invest in better explicit output-side LLM+MT cascades.
@@
We see that for both models, mean intermediate and final accuracy as well as mean are highest for Spanish, then Hindi, then Telugu (Table 1). It is unsurprising that intermediate accuracy and therefore final accuracy depend on the source language: task-solving success is naturally conditioned on comprehension of the task inputs. These findings justify exploring an input-side MT+LLM cascade for different language task inputs even for supported languages.
This also explains why is lowest for Telugu: with low intermediate accuracy, translation loss can only occur on very few inputs as a percentage of the total failure, which is dominated by failed task-solving cases. For example, since task-solving fails for of total inputs for Telugu-Cebuano with aya-23, task-solving failure dominates total failure even though the translation stage also fails on nearly all correct intermediate outputs.
4.4 Case study: Effect of scale
We repeat our experiments for language pairs on Llama-3.1-70B-Instruct and show the effect of using a larger model on in Figure 6. We find that the model shows similar intermediate accuracy on average as the model, but consistently higher final accuracy, or capability to produce the target language, resulting in lower . However, absolute values remain high. These preliminary results indicate that the translation barrier may continue to be a problem at scale.
4.5 Case study: Arithmetic Task
We conduct a small case study to investigate the translation barrier on an arithmetic task. We used word arithmetic English questions (such as “Eight times two equals?”), with numeral word answers, and prompted the model to respond in a target language, for variously resourced languages. We manually examine the outputs and observe the task-solvingtranslation pipeline, with the correct solved answers in HRLs in intermediate layers, with a switch to the target language in late layers. (See Appendix K for examples.) We report intermediate and final accuracy, and in Figure 7. We observe high especially for lower-resourced languages; these results are consistent with our findings on word translation. We invite work in better understanding the translation barrier for other tasks and model sizes.
@@ 5 Related Work
While multilingual LLMs are trained to perform well for several languages (Lin et al., 2022; Workshop et al., 2023; Aryabumi et al., 2024), their pre-training data is dominated by English (Xue et al., 2021; Chung et al., 2023; Li et al., 2025a). Etxaniz et al. (2024) and Zhao et al. (2024) show that explicit cascading with MT often outperforms end-to-end multilingual inference. Further, works such as Wendler et al. (2024) and (Schut et al., 2025) apply logit lens to intermediate layers and observe that multilingual LLMs “think” in English, or that intermediate representation spaces of these models lie close to English. This is reinforced by works studying neuron activations for multilingual processing (Foroutan et al., 2022; Tang et al., 2024; Kojima et al., 2024), which find language-specific neurons only in later model layers. Concurrent works to ours such as Lu et al. (2025) and Wang et al. (2025) demonstrate that late layer translation failure affects cross-lingual consistency and accuracy for factuality recall. Our work contributes to this discourse by formalizing and quantifying the translation barrier for a core generation task over a range of languages, with a focus on its impact on low-resource languages.
@@ 6 Conclusion
We formalize the translation barrier hypothesis for poor quality multilingual generation with LLMs. Working with a minimal generation task for language pairs, we visualize a task-solvingtranslation pipeline for multilingual generation, demonstrating its failure for low-resource languages. We then quantify the role of the translation barrier as an important culprit for low final performance, discussing its relevance for differently resourced source and target languages. Our findings contribute insights for future strategies in multilingual generation with LLMs.
@@
This work only explores the translation barrier hypothesis for a word translation task. It may be more difficult to assess task-solving success for more complex sequence-level tasks such as sentence-level machine translation, summarization, or open-ended instruction following in a principled manner using artifacts such as intermediate layer representations, since the nature of correct intermediate answer concepts could be less clear to ascertain. There may also be additional complexities to take into account when extending our setup to these tasks: for example, whether task-solving occurs “globally” over many parts or subtasks of a single input or “locally” for different fragments of the input, or the manner in which the target language syntax interacts with the ordering of the concepts in the task-solving stage. Thus, our hypothesis may be difficult to verify for other tasks.
We clarify that our findings regarding the role played by the translation barrier may be dependent on the nature and complexity of the task, as well as the usage of strategies such as task- or language-specific finetuning or inference-time techniques such as in-context learning to boost multilingual performance.
While we do not claim that the statement of the hypothesis holds for all tasks and settings, we believe that the line of investigation in our work has important implications for future directions to improve the multilingual capabilities in LLMs, especially regarding the potential of end-to-end versus cascaded or modular approaches for the same. Our goal is to invite discussion and further investigation of the translation barrier, as well as the most viable strategies for its mitigation.
@@
Finally, we use English as our prompt language regardless of the source language (i.e. the language of the term to be translated); this is as per work that says that this performs better (Dey et al., 2024). Testing the effect of different prompt languages may add further nuance to our findings.
@@
We do not anticipate any negative consequences of this work, which is intended to diagnose poor multilingual capabilities in large language models. We acknowledge that different language communities may have various and diverse needs from language technologies that may diverge from the points of focus of mainstream NLP, and not necessarily include LLM technologies (Bird, 2020). All the models we used are publicly available for research purposes: Aya-23-8B has a CC-BY-NC license, and the Llama family of models have a Llama 3.1 Community License Agreement, permitting research use. We will release all the artifacts created with the non-anonymized version of this paper. We used AI agents only for coding assistance (e.g. GitHub Copilot).